# On the Utility of Accounting for Human Beliefs about AI Behavior in Human-AI Collaboration

### v1

@@ 3.3 Designing Collaborative AI Agents
Collaborative AI agents. Using the methodology above, we have designed several AI agents based on different assumptions of the behavior of human counterparts.
@@
    Incorporating models of human behavior and beliefs. Finally we also train an AI agent that incorporates both the models of human behavior and beliefs into the design of AI agents.
The details of the setup and implementations are included in the appendix.
4 Experiments
@@
To evaluate our approaches, we have conducted three sets of experiments, consisting of both simulations and human-subject experiments. For our human-subject experiments, we have recruited in total $1,690$ participants from Amazon Mechanical Turk (MTurk) for multiple sets of experiments. The experiments are approved by the IRB of our institution. Workers were paid \$1 base payments with the potential for bonus payments in some experiments. The average hourly rate is approximately \$$\$14$ across all our experiments.
4.1 Experiment Environment: Grid World with Two Players and Multiple Goals
Our experiments are conducted in a grid world environment with two players and multiple goals. We have conducted two variations of the experiments. In the first variation, we designed the environments such that the two players are not playing in the same grid world. However, each of them has access to the full information of the environment and the actions of the other player. This variation abstracts away the interdependency of agent actions, meaning the agents’ actions do not influence each other directly. This allows us to focus on how humans and AI reason about each other’s goals and intentions. Note that we obtained qualitatively similar results in this first variation and the second, more complicated variation of environments described next. For simplicity, we have included the setup and results of the first variation in the appendix. Below, we describe the experiment environment and results of our second variation.
@@ 4.1 Experiment Environment: Grid World with Two Players and Multiple Goals
More specifically, as shown in Figure 2, in our second variation of the environment, the grid world is 8 by 8 in size, containing the positions of both players and four possible goals. The players can choose to move {Up, Down, Right, Left} or stay in the current grid. They can see each other’s positions and take actions simultaneously. Each player can navigate to two of the four goals. In particular, the human player can only reach one of the two goals colored blue, and the AI player can only reach one of the goals colored green. When both agents reach the same type of goal (both reaching a "star" or both reaching a "triangle"), they earn positive points. However, they will not earn points if they reach different goals or if they collide into each other (move into the same position). We set the maximum number of actions to be 20.
(a) The interface for experiment 1 & 3.
(b) The interface for experiment 2.
Figure 2: The interface for our human-subject experiment. In experiment 1, each participant is playing with an AI agent. The participants in experiment 1 are told what their goal is and only need to focus on reaching the goal without colliding with the AI agent. In experiment 2, each participant is provided traces of the behavior by other agents and is asked to infer which goal one agent is trying to reach. Finally, in experiment 3, the participants are not told which goal to reach, and they need to make decisions based on their beliefs over AI behavior. In experiments 1 and 3, the participants can only receive bonus rewards by reaching the same type of goals (star or triangle) as the AI agent.
In the first experiment, we recruit real-world humans to play against another pre-defined AI agent. We then examine the effectiveness of utilizing behavioral cloning as a model to mimic human behavior. Note that using behavioral cloning to model human behavior is extensively adopted in the literature, and the purpose of the first experiment serves as the foundation to develop our human belief models and the design of collaborative AI agents.
To evaluate our proposed approaches in modeling human beliefs and designing cooperative AI, we have designed and conducted additional two sets of experiments. In the second set of experiments, we provide participants with different traces of actions from other agents, and ask participants to infer the goal of the agents. This experiment helps us evaluate whether our belief model leads to better predictions of human beliefs over others’ behavior. Afterwards, in the third experiment, we conduct experiments similar to the first experiment. We team up AI agents and human players to examine the team performance of our design of collaborative AI when working with real-world human participants. But unlike the first experiment, we don’t reveal any information about goals the players are trying to reach.
4.2 Experiment 1: Training and Evaluating Human Behavior Models
@@
Experiment setup. We recruited 190 workers from Amazon Mechanical Turk. Each recruited worker was asked to play 30 navigation games within a grid world with a pre-defined AI model, as shown in Figure 2a. To collect diverse data, we designed three AI models to play with humans: random AI (selecting a random goal and taking the shortest path), self-play AI (trained using self-play), and optimal AI (optimized via joint optimization).
Note that in this experiment, our goal is to develop a model of human behavior that outputs a sequence of actions given a goal. Therefore, in experiment 1, we provided additional instructions to participants about which goal they should reach in each round, and the AI models controlled the other player to move towards the same type of goals. In our later experiment 3, we did not provide these instructions, and humans needed to infer the goal of the AI agent.
Evaluation of data-driven behavior models. We divided the collected data of human actions into three sets: training, validation, and testing. The training set comprised data from 152 workers, including approximately 136,000 instances of user decisions, while the validation and testing sets each contained data from 19 workers, amounting to around 17,000 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model, where the input is the current environment layout, and the output predicts the next human action. We fine-tuned the hyperparameters, such as learning rate, hidden layer size, and L2 penalty, based on validation errors.
Our results suggest that human behavior is noisy and is significantly away from being optimal. Even when human participants are provided with suggested goals during data collection, there is only a $55.0\%$ chance that both players will reach the same type of goals across all treatments. Besides, there is a notable chance of two players colliding (about $15.8\%$) or ending in different types of goals (about $25.6\%$). We compared the performance of our learned model to a model predicated on optimal agent behavior, defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly reveal that human behavior deviates significantly from the assumed optimality. This deviation highlights the importance of incorporating a realistic model of human behavior into human-AI cooperation frameworks.
Table 1: The prediction accuracy for human behavior assuming optimal behavior and using data-driven model in Experiment 1.
@@
4.3 Experiment 2: Evaluating Human Belief Models
We next examine our model of human beliefs. Specifically, we investigate whether we can design AI behavior such that it is easier for humans to infer the goal of the AI agent.
@@ 4.3 Experiment 2: Evaluating Human Belief Models
In our experiments with a simplified environment, as included in the appendix, we directly examined how accurately our belief model can infer human beliefs by comparing the predictions of our models with the direct solicitation of users’ inferences about others’ behavior. The results suggest that our model is more accurate in predicting human beliefs compared to baselines. However, human beliefs are generally very noisy, and if we randomly draw a behavior trace from another agent and ask humans to predict the goal, human prediction accuracy is close to random guessing. Therefore, we next shift our focus to whether AI agents can design their action plans to make it easier for humans to infer the goal. In this more complicated environment, we have directly examined this follow-up question of designing AI policies that are explicable to humans.
Experiment setup. The experiment setup is presented in Figure 2b. In addition to player positions and goal positions, we also display behavioral traces, which are sequences of actions taken by previous players. For each participant, we show them the traces of two players and ask the participant to infer which one of the two goals those agents are trying to reach.
We compare two belief models. The first one is the standard level-1 model: humans assume the other agent is a level-0 agent that takes the optimal decision (i.e., taking the shortest path towards the goal). The second one is our proposed behavioral level-1 model: humans assume the other agent is also a suboptimal decision-maker, taking actions in the same way as themselves. In our implementation, we leverage the data-driven behavioral model derived from Experiment 1 as the user behavioral model. More concretely, we construct a belief model using Bayesian inference, as described in Section 3.2, using the corresponding human model of behavior.
Explicable AI policy. As mentioned earlier, humans generally struggle to infer the goals of other agents from behavioral traces randomly drawn from history (see the appendix for detailed results). Therefore, in this experiment, instead of displaying a randomly drawn historical behavioral trace, we display the behavioral trace of an AI agent that aims to make its behavior explicable. In particular, based on the developed belief models, we train AI agents to not only achieve their goals but also to maximize the likelihood that humans can accurately infer these goals from their behavior. Our implementation utilizes self-play and awards an additional bonus when the goal inference aligns with the predictions of a human belief model, in addition to the rewards for goal achievement. The bonus is proportional to the log likelihood of belief model inference $log(\pi(a|s,g))$. By conducting experiments with this explicable AI policy, we simultaneously evaluate whether our belief models are accurate and whether our design of explicable AI policy is effective.
Figure 3: The simulation results in Experiment 2.
Simulations. We design the AI agents to incorporate the belief models of standard level-1 agents and behavioral level-1 agents, enabling them to adopt policies that simplify the task of goal inference for humans. For each environment with a given goal, we generated the behavioral traces of both AI policies. The goal is to examine whether the explicable AI with our belief model makes it easier for humans to infer the goal of the AI. We first conducted simulations assuming humans are behavioral level-1 agent. The results are shown in Figure 3. The explicable AI equipped with the belief model of behavioral level-1 agents leads to behavioral traces that are easier to infer. Note that this result is not surprising is not surprising since we assume humans infer the goal following the beliefs of behavioral level-1 agent. The results provide evidence that our design of explicable AI is effective.
Figure 4: The results of human-subject experiments in Experiment 2.
Human-subject experiments. We next conducted human-subject experiments to assess whether our approach results in actions that make it easier for real humans to infer the goal. We recruited 200 workers from Amazon Mechanical Turk, randomly assigning them to one of two treatments. Each participant was tasked with identifying the goals of the player in 30 different scenarios. The length of actions is drawn from the range 4 to 8. To incentivize effort, participants were awarded a \$$\$0.03$ bonus for each correctly identified goal. The results, as shown in Figure 4, demonstrate that humans are better at inferring the goal of the explicable AI agent, that assumes humans are behavioral level-1 agents. The difference is statistical significant with $p>0.0001$. This comparison demonstrates that when coupled with the more accurate human belief models, we can indeed design AI policy to be explicable, i.e., making it easier for humans to infer AI goals from observing AI actions.
4.4 Experiment 3: Evaluating the Design of Collaborative AI Agents
We now examine our design of collaborative AI agents. As described in Section 3.3, we train collaborative AI agents using three different corresponding human models and utilize both simulations and human-subject experiments to examine the effectiveness of our design.
Experiment setup. The setup of our Experiment 3 is the same as Experiment 1, shown in Figure 2a. But human participants will not receive any hint about which goal their teammate or themselves are suggested to reach.
Simulations. We first run simulations to evaluate the performance of human-AI teams with different design of collaborative AI. The evaluation is based on $10,000$ randomly generated environments. However, we further filter out cases where the distance between the two goals for the same player is smaller than 3 to encourage models to adjust their behavior based on the inference of their teammate’s actions. We examine the performance of three designs of collaborative AI using the methodology described in Section 3.3:
@@ 4.4 Experiment 3: Evaluating the Design of Collaborative AI Agents
    Self-Play is the AI that is trained assuming they are playing with itself.
@@
    Behavior-AI is the AI that assumes the human partner is the behavioral level-00 agent.
@@
    Belief-AI is the AI that assumes the human partner is the behavioral level-$1$ agent.
We partner the collaborative AI agents with the three simulated human agents and measure the collaborative performance. The simulation results, shown in Table 2, highlight the importance of selecting appropriate human models for training collaborative AI. When an AI agent is paired with the human model that is used to train the AI, the collaborative performance is better compared to pairing with other models.
Table 2: Simulation results of collaborative performance over 10k testing cases in experiment 3. Column players are different AI agents, and row players are different simulated human models.
@@
Self-play Behavior-AI Belief-AI
Self-play AI 0.6245 0.5250 0.6177
Behavior Model 0.4902 0.6334 0.5755
Behavior&Belief 0.6411 0.6574 0.7675
Figure 5: The average human-AI performance in our human-subject experiment in Experiment 3.
@@
Human-subject experiments. To examine the performance of our collaborative AI design when pairing with real humans, we recruited 200 workers and divided them into three treatment groups, each interacting with a different design of AI. Participants were tasked with playing 30 games, preceded by three tutorial games designed to familiarize them with the gameplay. Participants can receive a \$0.05 bonus payment for reaching the same goal with the AI agent for each of the 30 tasks.
@@
Figure 5 presents the average collaborative reward, highlighting that AI trained with models of human behavior and beliefs achieved the highest collaborative performance when working with real humans. This outcome not only validates our design of AI but also suggests that real humans’ actions align with the behavior and belief models in this environment. Statistical analysis shows significant differences in the performance of AI models paired with humans, with $p<0.0001$ for comparisons between self-play and AI accounting for human behaviors, and a $p$ value of $0.0014$ for comparisons between AI accounting for human behaviors versus accounting for both behavior and belief. These results demonstrate that incorporating human beliefs into the deign of AI agents enhances collaborative performance when working with real humans.
5 Conclusion

### v2

@@ 3.4. Designing Collaborative AI Agents
Collaborative AI agents. Using the methodology above, we have designed several AI agents based on different assumptions of the human counterpart.
@@
    Incorporating models of human beliefs. Finally we also train an AI agent that incorporates both the models of human behavior and beliefs into the design of AI agents.
In addition to the AI agents trained using our methodologies, we also include two state-of-the-art baselines, Fictitious Co-Play (Strouse et al., 2021) and Maximum Entropy Population-based Training (Zhao et al., 2023), for comparison in our experiments. Both algorithms develop a single best response model in response to a population agents with diverse behaviors, thereby enhancing the robustness and generality of the learned best response model. The details of the setup and implementations are provided in the appendix.
4. Experiments
@@
To evaluate our approaches, we have conducted multiple sets of human-subject experiments through recruiting close to 1,000 participants from Amazon Mechanical Turk (MTurk). The experiments are approved by the IRB of our institution. Workers were paid \$1 base payments with the potential for bonus payments in some experiments. The average hourly rate is approximately \$$\$14$ across all our experiments.
4.1. Experiments Overview
The main purpose of our experiments is to demonstrate that incorporating human beliefs about AI intentions into the design of AI agents can significantly improve the performance of human-AI collaboration. Specifically, we conducted three sets of experiments, beginning with examining the efficacy of our models of human behavior and beliefs, and progressing to evaluate the performance improvements.
In the first experiment, we examine the effectiveness of using behavioral cloning as a model to mimic human behavior. We recruit real-world participants to play against a predefined AI agent to collect their behavioral traces and then evaluate the prediction accuracy of using behavioral cloning to model human behavior. In the second set of experiments, we examine whether we can leverage the model of human beliefs to design an explicable AI policy (as described in Section 3.3). Specifically, we recruit real-world participants, provide them with behavioral traces generated by different AI policies, and evaluate whether the AI policies that account for human belief models indeed produce behavioral traces that make it easier for humans to infer AI intentions. Finally, in the third experiment, we examine whether the collaborative AI that accounts for human beliefs lead to higher performance in human-AI collaboration. Again we recruit real-world human participants, pair them with different AI agents (as described in Section 3.4), and measure the overall team performance.
4.2. Experiment Environment
@@
Our experiments are conducted in a grid world environment with two players and multiple goals. [^3] We conducted experiments in two variations of the environment. In the first variation, we designed the environments such that the two players are not in the same grid world. However, each player has access to full information about the environment and the actions of the other player. This variation abstracts away the interdependency of agent actions, meaning the agents’ actions do not directly influence each other. This allows us to focus on how humans and AI reason about each other’s goals and intentions. In the second variation, we consider a more complex and realistic situation where both players are in the same grid world environment and may collide with each other. We obtained qualitatively similar results in both variations. Due to space constraints, we have included the setup and results of the second variation in the main paper and moved the results of the first variation to the appendix. More specifically, as shown in Figure 2, the grid world is 8 by 8 in size and contains the positions of both players and four possible goals. The players can choose to move {Up, Down, Right, Left} or stay in their current grid. They can see each other’s positions and take actions simultaneously. Each player can navigate to two of the four goals. In particular, the human player can only reach one of the two goals colored blue, and the AI player can only reach one of the goals colored green. When both agents reach the same type of goal (e.g., both reaching a ”star” or both reaching a ”triangle”), they earn positive points. However, they do not earn points if they reach different types of goals or if they collide with each other (move into the same position). We set the maximum number of actions to be 20.
Figure 2. The decision-making environment in simulations and human-subject experiments. The human-AI team will receive a reward by reaching the same type of goals (star or triangle) within the time limit. They will receive no reward if they collide with each other or reach different types of goals.
4.3. Experiment 1: Evaluating Behavior Models
@@
Experiment setup. We recruited 190 workers from Amazon Mechanical Turk. Each recruited worker was asked to play 30 navigation games within a grid world with a predefined AI model, as shown in Figure 2. To collect diverse data, we randomly paired recruited workers with one of three AI models: a random AI (selecting a random goal and taking the shortest path), a self-play AI (trained using self-play), and an optimal AI (optimized via joint optimization).
In this experiment, our goal is to develop a model of human behavior that outputs a sequence of actions given a goal. Therefore, in Experiment 1, we provided participants with additional instructions about which goal they should reach in each round, and the AI models controlled the other player to move towards the same type of goals. In our later Experiment 3, we did not provide these instructions, and humans needed to infer the goal of the AI agent.
Experiment results. We divided the collected data of human actions into three sets: training, validation, and testing. The training set comprised data from 152 workers, including approximately 136,000 instances of user decisions, while the validation and testing sets each contained data from 19 workers, amounting to around 17,000 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model, where the input is the current environment layout, and the output predicts the next human action. We fine-tuned the hyperparameters, such as learning rate, hidden layer size, and L2 penalty, based on validation errors.
Table 1. The prediction accuracy for human behavior was evaluated assuming optimal behavior and using a data-driven model in Experiment 1. The results show that human behavior deviates significantly from optimality, and our data-driven model provides a much more accurate prediction of human behavior.
@@
Our results suggest that human behavior is significantly away from being optimal. Even when human participants are provided with suggested goals during data collection, there is only a $55.0\%$ chance that both players will reach the same type of goals across all treatments. Besides, there is a notable chance of two players colliding (about $15.8\%$) or ending in different types of goals (about $25.6\%$). We compared the performance of our learned model to a model predicated on optimal agent behavior, defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly reveal that human behavior deviates significantly from the assumed optimality. This deviation highlights the importance of incorporating a realistic model of human behavior into human-AI cooperation frameworks.
4.4. Experiment 2: Evaluating Belief Models
@@
We next evaluate our model of human beliefs. In particular, we examine whether an AI policy that incorporates human behavior models can produce actions that are more explicable to humans, i.e., easier for humans to infer AI intentions based on AI actions. [^4] We have also directly tested the accuracy of our belief model in inferring human beliefs by comparing its predictions with direct solicitations of users’ inferences about others’ behavior. The results, included in Appendix A.3, suggest that our model is more accurate in predicting human beliefs compared to the baselines. However, they also demonstrate that human beliefs are generally quite noisy—if we randomly draw a behavior trace from another agent and ask humans to predict the goal, human prediction accuracy is close to random guessing. Therefore, we shift our focus in this experiment to whether AI agents can design their action plans to make it easier for humans to infer the goal.
Experiment setup. In this experiment, each recruited participant is presented with the behavioral traces from both agents in the environment, as shown in Figure 3. Each behavioral trace is a sequence of actions generated by an AI policy for a player. Participants are then asked to infer the intention of each agent, i.e., the goal each agent is trying to reach. If participants can better infer the goal from the behavior generated by a certain AI policy, it indicates that the AI policy is more explicable in that it generates actions that make it easier for humans to infer AI intentions.
Figure 3. The experiment interface in Experiment 2. Participants are given the behavioral traces of the agents and are asked to infer the goal each agent aims to reach.
AI policy. We now describe the set of AI policies that are used for the evaluations. As baselines, the first two AI policies are not intended to be explicable.
@@
    Inexplicable-Optimal-L0: This AI policy that takes the shortest path towards the goal.
@@
    Inexplicable-Behavioral-L0: This AI policy takes actions that follow the data-driven behavioral model learned in experiment 1.
We then include two AI policies that take actions intended to be explicable. Note that to design AI policies to be explicable, we need to account for human belief models in the design of AI policies, as the AI will try to take actions that make it easier for humans to infer the goal, assuming models of human beliefs that describe how humans make inferences. The two belief models we account for are: (1) the classical level-$1$ model, where the human assumes the other agent is taking optimal level-00 actions, and (2) the behavioral level-$1$ model, where the human assumes the other agent is taking behavioral level-00 actions as described in Section 3.2. Coupled with these two belief models, and using the methodology described in Section 3.3, we can train the following two explicable AI policies.
@@
    Explicable-Optimal-L1: This AI policy is trained to be explicable but assumes that humans belief model is the classical level-$1$ model, where they assume the other agent is optimal.
  • $\bullet$
    Explicable-Behavioral-L1: This AI policy is trained to be explicable but assumes that human belief model is the behavioral level-$1$ model, where they assume the other agent is behavioral level-00.
Figure 4. The results of the human-subject experiments in Experiment 2 show that when using our approach to design explicable AI, the accuracy for humans to infer AI intentions is higher than with other AI policies. Moreover, when coupled with our belief model, it results in an AI policy that makes it the easiest for humans to infer AI intentions.
Experiment results. We recruited 300 workers from Amazon Mechanical Turk, randomly assigning them to one of four treatments, where each treatment represents showing participants traces generated by one of the AI policies above. Each participant was tasked with identifying the goals of the player in 30 different scenarios. The length of actions is drawn from the range 4 to 8. To incentivize effort, participants were awarded a \$$\$0.03$ bonus for each correctly identified goal.
The experiment results are shown in Figure 4. First, we observe that when the AI is not attempting to be explicable, humans generally struggle to infer AI intentions, with the accuracy of correctly inferring the goal for both inexplicable policies being around 55% to 60%, only slightly better than random guessing. However, if we use our methodology from Section 3.3 to train the AI policy to be explicable, it significantly improves the accuracy of humans in inferring AI intentions. Moreover, when the training is coupled with our belief models, the accuracy for humans to infer AI intentions is the highest, exceeding 70%.
4.5. Experiment 3: Evaluating AI Design
We now examine our design of collaborative AI agents. Our goal is to demonstrate that the collaborative AI design that accounts for human beliefs could lead to improved human-AI collaborative performance. We utilize both simulations and human-subject experiments to examine the effectiveness of our design.
Table 2. Simulation results of collaborative performance over 10,000 testing cases in Experiment 3. The column players represent different AI agents, and the row players represent different simulated human models. The results demonstrate that while the zero-shot methods (FCP and MEP) show consistent and stable performance, when human agents possess belief models and infer AI intentions to act accordingly, designing AI that accounts for human beliefs brings significant performance improvements.
@@ 4.5. Experiment 3: Evaluating AI Design
Self-Play FCP MEP Behavior-AI Belief-AI
Self-play 0.6245 0.6136 0.6131 0.5250 0.6177
Behavior Model 0.4902 0.5782 0.5927 0.6334 0.5755
Behavior&Belief 0.6411 0.6255 0.6484 0.6574 0.7675
Experiment setup. The setup of our Experiment 3 is the same as Experiment 1, shown in Figure 2. But human participants will not receive any hint about which goal their teammate or themselves are suggested to reach.
Collaborative AI design. As described in Section 3.4, we train collaborative AI agents using three different corresponding human models.
  • $\bullet$
    Self-Play: the collaborative AI agent that is trained assuming they are playing with itself.
  • $\bullet$
    Behavior-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-00 agent.
  • $\bullet$
    Belief-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-$1$ agent, i.e., it accounts for human belief models into the design of collaborative AI.
In addition to the three collaborative AI agents trained using our methodology, we also include two state-of-the-art baselines for comparison. These methods leverage zero-shot population-based training (PBT) techniques. While they do not explicitly account for specific models of human behavior, they are designed to be robust in collaborating with different teammate behaviors.
  • $\bullet$
    FCP: Fictitious Co-Play is a variant of fictitious play proposed by Strouse et al. (2021). It consists of two stages: in the first stage, a population of agents is trained independently using different random seeds and checkpoints; in the second stage, a model is trained as the best response to the agent population. In our experiment, we use the model trained with 8 partners.
  • $\bullet$
    MEP: Maximum Entropy Population-based training is proposed by  Zhao et al. (2023). The agent population is trained using population entropy bonus to increase the diversity of agents, then a single best response model is trained with respect to the agent population. In our experiment, we use the model trained with 8 partners.
Simulations. We first run simulations to evaluate the performance of different pairings of human models and collaborative AI designs, where we assume the human agent is optimal (self-play), an behavioral level-00 agent (Behavioral Model), and behavioral level-$1$ agent (Behavior&Belief). The evaluation is based on $10,000$ randomly generated environments which filter out cases where the distance between the two goals for the same player is smaller than 3 to allow models to adjust their behavior based on the inference of their teammate’s actions.
We partner the collaborative AI agents with the three simulated human agents and measure the collaborative performance. The simulation results are shown in Table 2. For both the FCP and MEP baselines, despite the absence of human data in their training processes, they outperform the self-play AI when paired with the human behavior model, demonstrating their adaptability to working with teammates exhibiting unknown behavior. However, we also observe that incorporating the appropriate human behavior model directly into the training process can further enhance human-AI collaborative performance, as demonstrated by the superior performance of Behavior-AI when paired with the behavior model. Finally, when working with a behavioral level-1 agent that attempts to infer AI intentions, our proposed Belief-AI stands out, significantly outperforming other AI agents, which perform similarly to each other.
These results highlight the importance of selecting appropriate human models for training collaborative AI. Furthermore, under conditions where humans attempt to infer AI intentions and act accordingly, designing AI that accounts for human beliefs can significantly improve human-AI collaborative performance.
@@
Human-subject experiments. While the simulation results are promising, the significant performance improvement from incorporating human beliefs about AI intentions into AI design is based on the assumption that humans do indeed infer AI intentions and act accordingly. To examine whether this performance improvement carries over to collaboration with real humans, we recruited 400 workers and randomly assigned them into five treatment groups, each interacting with a different AI design. Participants were tasked with playing 30 games, preceded by three tutorial games designed to familiarize them with the gameplay. Participants could receive a \$0.05 bonus payment for reaching the same type of goal as the AI agent in each of the 30 tasks.
@@
The experiment results are shown in Figure 5. They demonstrate that the collaborative AI trained with models of human behavior and beliefs achieved the highest collaborative performance when working with real humans. This outcome not only validates our AI design but also suggests that real human actions align with the behavior and belief models in this environment. Statistical analysis shows significant differences in the performance of AI models paired with humans, with $p<0.0001$ for comparisons between self-play AI and AI accounting for human behavior, and $p<0.0001$ for comparisons between AI accounting for human behavior versus accounting for both behavior and beliefs. The differences between MEP and FCP, as well as between FCP and behavioral-AI, are not significant ($p>0.1$). The difference between self-play AI and any other treatment is significant ($p<0.0001$). These results demonstrate that incorporating human beliefs into the design of AI agents enhances collaborative performance when working with real humans.
Figure 5. The human-AI collaborative performance in our human-subject experiment in Experiment 3 shows that designing collaborative AI that accounts for human beliefs (Belief-AI treatment) significantly outperforms all other baselines when working with real humans.
5. Conclusion